Cost 150 130 120 150 AIS 1,000| $ 130 130 210 910 1,110 Hint#1
Answer #1
Cases Sales revenue Beginning inventory Purchases Total available Ending inventory Cost of goods sold Gross Profit Expenses Pretax income(loss)
A 1000 300 850 1150 420 730 270 150 120
B 1160 130 980 1110 300 810 350 130 220
C 640 130 460 590 230 360 280 120 160
D 840 150 590 740 210 530 310 150 160
E 1060 200 910 1110 510 600 460 520 -60
Case A:
Total available 1150 300+850
Cost of goods sold 730 =1150-420
Gross Profit 270 =1000-730
Pretax income(loss) 120 =270-150
Case B:
Total available 1110 =130+980
Ending inventory 300 =1110-810
Cost of goods sold 810 =1160-350
Gross Profit 350 =130+220
Case C:
Sales revenue 640 =360+280
Purchases 460 =590-130
Total available 590 =360+230
Pretax income(loss) 160 =280-120
Case D:
Beginning inventory 150 =740-590
Total available 740 =530+210
Cost of goods sold 530 =840-310
Gross Profit 310 =150+160
Case E:
Beginning inventory 200 =1110-910
Ending inventory 510 =1110-600
Cost of goods sold 600 =1060-460
Expenses 520 =460+60